OpenStudy (anonymous):

(5e^5x) + 11 = 3x+41

OpenStudy (gorv):

what we need to do ..???

OpenStudy (gorv):

@Jared25

OpenStudy (gorv):

subtract 11 from both side

OpenStudy (gorv):

5e^5x+11-11=x+43-11

OpenStudy (gorv):

what u will get ??

OpenStudy (anonymous):

5e^5x=3x+30

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

find approximate solutions of the equation

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Just plug in into the calculator and you will get your answer.

OpenStudy (anonymous):

I tried it didnt work

OpenStudy (anonymous):

Answer should be 0.163... Just recheck it using your calculator.

OpenStudy (anonymous):

thats what i got too but it should be 0.365